OpenStudy (anonymous):

Brett deposited $3200 into a savings account for which interest is compounded quarterly at a rate of 2.4%. How much interest will he earn after 15 years?

OpenStudy (johnweldon1993):

\[\large A = P(1 + \frac{r}{n})^{nt}\] P = 3200 r = .024 t = 15 n = quarterly = 4 times a year = 4 so we have \[\large A = 3200(1 + \frac{.024}{4})^{4 \times 15}\] So what does A equal?

OpenStudy (anonymous):

4576

OpenStudy (johnweldon1993):

Hmm I receive a result of 4581.7 So whatever result you get from that equation....you take your original 3200 from it.. what we calculated was what the amount would be after the 15 years....and that is the 4581.7 Since we what to know how much interest that was....we take the new amount...and subtract the original amount... so \[\large 4581.7 - 3200 = ?\]

OpenStudy (anonymous):

1381.7

OpenStudy (johnweldon1993):

That would be correct